Versions Compared


  • This line was added.
  • This line was removed.
  • Formatting was changed.


Chris and I made a trip to MFO to move the 7m Li-COR up (sn 0813) up to 15m and to remove the dead instrument from 15m.  The tower is therefore operating again with three levels of water vapor and CO2 fluxes (2m, 16m, 43m).  See Chris' entry for details of this instrument swap and timing.